Developing an Equation from Average Costs The America Dog and Cat Hotel is a pet hotel located in Las Vegas. Assume that in March, when dog-days (occupancy) were at an annual low of 500, the average cost per dog-day was $14. In July, when dog-days were at a capacity level of 4,000, the average cost per dog-day was $7. (a) Develop an equation for monthly operating costs. (Let X = dog-days per month) Total cost =Answer + Answer * X

(b) Determine the average cost per dog-day at an annual volume of 24,000 dog-days.

$Answer
Answer #1

ANSWER :-

  • The problem can be addressed by the use of High Low Method to find Variable cost per dog days (units) and Fixed cost.
  • Workings

Units

Cost

High Level

4,000

$28,000.00 [4000 x $ 7]

Low Level

500

$7,000.00 [ 500 x $ 14]

Difference

3,500

$21,000.00

A

Difference in Cost

$21,000.00

B

Difference in units

3,500

C = A/B

Variable cost per unit

$6.00

Working

High Level

Low Level

A

Total Cost

$28,000.00

$7,000.00

B

Total Units

4000

500

C

Variable cost per unit

$6.00

$6.00

D = B x C

Total Variable cost

$24,000.00

$3,000.00

E = A - D

Total Fixed Cost

$4,000.00

$4,000.00

==> Requirement [a]

Variable cost = $ 6 per dog day
Fixed Cost = $ 4000

Equation:
Total Cost = $ 4,000 + $ 6 * X

==> Requirement [b]

Total Cost = $ 4,000 + $ 6 * X

Total Cost = 4000 + ($6 x 24000 dog days)
= 4000 + 144000

= $ 148,000

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************